# On the Inductive Biases of Demographic Parity-based Fair Learning Algorithms

### v1

@@ Abstract
Fair supervised learning algorithms assigning labels with little dependence on a sensitive attribute have attracted great attention in the machine learning community. While the demographic parity (DP) notion has been frequently used to measure a model’s fairness in training fair classifiers, several studies in the literature suggest potential impacts of enforcing DP in fair learning algorithms. In this work, we analytically study the effect of standard DP-based regularization methods on the conditional distribution of the predicted label given the sensitive attribute. Our analysis shows that an imbalanced training dataset with a non-uniform distribution of the sensitive attribute could lead to a classification rule biased toward the sensitive attribute outcome holding the majority of training data. To control such inductive biases in DP-based fair learning, we propose a sensitive attribute-based distributionally robust optimization (SA-DRO) method improving robustness against the marginal distribution of the sensitive attribute. Finally, we present several numerical results on the application of DP-based learning methods to standard centralized and distributed learning problems. The empirical findings support our theoretical results on the inductive biases in DP-based fair learning algorithms and the debiasing effects of the proposed SA-DRO method.
@@ 1 Introduction
A widely-used approach to fair supervised learning is to include a fairness regularization penalty term in the learning objective that quantifies the level of fairness violation according to a fairness notion. A standard fairness notion is the demographic parity (DP) aiming toward a statistically independent prediction variable $\widehat{Y}$ of a sensitive attribute $S$. Therefore, a DP-based fairness regularization metric should be a measure of the dependence of the prediction $\widehat{Y}$ on the sensitive attribute $S$. In the literature, several dependence measures from statistics and information theory have been attempted to develop DP-based fair learning methodologies.
In practice, the applications of standard DP-based fair classification methods usually succeed in significantly reducing the DP fairness violation, while the model’s original accuracy on test data can be mostly preserved. Therefore, an accuracy-based evaluation of the DP-based trained models often suggests that the improvement in the DP fairness metric can be significantly higher than the loss in the model’s prediction accuracy. On the other hand, well-known studies including [1] and [2] have raised concerns about the potential impacts of DP-based fairness evaluation, which can disproportionately increase the inaccuracy rate among minority subgroups. To address the concerns, [2] propose and promote a different fairness notion, equalized odds (EO), where the goal is a prediction variable $\widehat{Y}$ that is conditionally independent of the sensitive attribute $S$ given the true label $Y$.
@@
We present the results of several numerical experiments on the potential biases of DDP-based fair classification methodologies to the sensitive attribute possessing the majority in the dataset. Our empirical findings are consistent with the theoretical results, suggesting the inductive biases of DP-based fair classification rules toward the sensitive attribute-based majority group. On the other hand, our results indicate that the DRO-SA-based fair learning method results in fair classification rules with a lower bias toward the label distribution under the majority sensitive attribute.
Furthermore, to show the impacts of such inductive biases in practice, we analyze the fair classification task in a federated learning context where multiple clients attempt to train a decentralized model. We focus on a setting with heterogeneous sensitive attribute distributions across clients where the clients’ majority sensitive attribute outcome may not agree. Figure 1 illustrates such a federated learning scenario over the Adult dataset, where Client 1’s majority sensitive attribute (female samples) is different from the network’s majority group (male samples), and consequently, Client 1’s test accuracy with a DP-based fair federated learning is significantly lower than the test accuracy of a localized fair model trained only on Client 1’s data. Such numerical results question the client’s incentive in participating in fair federated learning. The following is a summary of this work’s main contributions:
@@ 2 Related Works
Fairness Violation Metrics. In this work, we focus on the learning frameworks aiming toward demographic parity (DP). Since enforcing DP to strictly hold could be costly and damaging to the learner’s performance, the machine learning literature has proposed applying several metrics assessing the dependence between random variables, including: the mutual information: [3, 4, 5, 6, 7], Pearson correlation [8, 9], kernel-based maximum mean discrepancy: [10], kernel density estimation of the difference of demographic parity (DDP) measures [11], the maximal correlation [12, 13, 14, 15], and the exponential Renyi mutual information [16]. In our analysis, we mostly focus on a DDP-based fair regularization scheme, while we show only weaker versions of the inductive biases could further hold in the case of mutual information and maximal correlation-based fair learning algorithms.
In addition to DP, the notions of equalized odds and equal opportunity [2] are standard fairness notions in the literature, where the learner aims for a conditionally independent decision variable $\widehat{Y}$ of sensitive attribute $Z$ given label $Y$. We note that the mentioned frameworks based on information-theoretic dependence measures can be utilized to target equalized odds, where the dependence measure should be conditioned to the label $Y$’s outcome. In most cases, fair classifiers based on one of the fairness notions are not guaranteed to achieve good performance on the others. Therefore, our theoretical and numerical findings do not apply to the equalized odds fairness notion and the extension of the statistical dependence measure-based learning algorithms targeting equalized odds.
Fair Classification Algorithms. Fair machine learning algorithms can be classified into three main categories: pre-processing, post-processing, and in-processing. Pre-processing algorithms [17, 18, 19] transform biased data features into a new space where labels and sensitive attributes are statistically independent. Post-processing methods such as [2, 20] aim to alleviate the discriminatory impact of a classifier by modifying its ultimate decision. The focus of our work focus is only on in-processing approaches regularizing the training process toward DP-based fair models. Also, [21, 22, 23] propose distributionally robust optimization (DRO) for fair classification; however, unlike our method, these works do not apply DRO on the sensitive attribute distribution to reduce the biases.
@@ 3.1 Fair Supervised Learning
In the supervised learning problem, the learner selects a function $f\in\mathcal{F}$ where $\mathcal{F}$ is the set of prediction functions mapping the observed $(\mathbf{X},S)$ to the label space $\mathcal{Y}$. We use loss function $\ell\mathrel{\mathop{:}}\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ to quantify the loss $\ell(y,\widehat{y})$ when predicting $\widehat{y}$ under a true label $y$. Specifically, we consider the 0/1 loss $\ell_{0/1}(\widehat{y},y)=\mathbf{1}(\widehat{y}\neq y)$, where $\mathbf{1}(\cdot)$ denotes the indicator function. The primary goal of the fair supervised learner is to find prediction rules $f\in\mathcal{F}$ achieving smaller values of risk function $\mathbb{E}_{(\mathbf{X},Y,S)\sim P}\bigl{[}\ell(f(\mathbf{X},S),Y)\bigr{]}$ while having little dependence on $Z$ according to the factors explained in the next subsections.
@@ 3.2 Fairness Criteria
  2. 2.
    Equalized Odds (EO) [2] is a fairness condition requiring the predicted label $Y$ to be conditionally independent from sensitive attribute $S$ given actual label $Y$, i.e. for every $s\in\mathcal{S},y,\widehat{y}\in\mathcal{Y}$
@@
    $\centering\mathrm{DEO}(\widehat{Y},S|Y)=\sum_{s\in\mathcal{S},y,\hat{y}\in\mathcal{Y}}\Bigl{|}\,\@add@centering$ $P(\widehat{Y}=\widehat{y}\,\big{|}\,Y=y,S=s)-P\bigl{(}\widehat{Y}=\widehat{y}\,\big{|}\,Y=y\bigr{)}\Bigr{|}$
@@ 3.3 Dependence Measures for Fair Supervised Learning
    • Mutual Information (MI): The mutual information $I(Y;S)$ is a standard measure of the dependence between random variables $Y$ and $S$, defined as
      $I(Y;S)\mathrel{\mathop{:}}=\sum_{y\in\mathcal{Y},s\in\mathcal{S}}P_{Y,S}(y,s)\log\frac{P_{Y,S}(y,s)}{P_{Y}(y)P_{S}(s)}$
@@
    • Maximal Correlation (MC): The maximal correlation $\rho_{m}(Y,S)$ is the maximum Pearson correlation $\rho_{P}\bigl{(}f(Y),g(S)\bigr{)}=\frac{\mathrm{Cov}(f(Y),g(S))}{\sqrt{\mathrm{Var}(f(Y))\mathrm{Var}(g(S))}}$ between $f(Y)$ and $g(S)$ over all functions $f,\,g$ with zero-mean and unit-variance $f(Y)$ and $g(S)$ random variables:
      $\rho_{m}(Y,S)\mathrel{\mathop{:}}=\max_{\begin{subarray}{c}f,g\mathrel{\mathop{:}}\;\mathbb{E}[f(Y)]=\mathbb{E}[g(S)]=0\\ \qquad\,\mathbb{E}[f^{2}(Y)]=\mathbb{E}[g^{2}(S)]=1\end{subarray}}\rho_{P}\bigl{(}f(Y),g(S)\bigr{)}$
    • Exponential Rényi Mutual Information (ERMI): The ERMI between random variables $Y$ and $S$ is
      $\rho_{E}(Y,S)\mathrel{\mathop{:}}=\chi^{2}(P_{Y,S};P_{Y}\times P_{S}),$
      i.e, the $\chi^{2}$-divergence between the joint distribution $P_{Y,S}$ and the product of marginal distributions $P_{Y}\times P_{S}$. Similar to KL-divergence, $\chi^{2}$-divergence is an $f$-divergence $d_{f}(P,Q)$ with $f(t)=(t-1)^{2}$. Also, $\rho_{E}(Y,S)$ is zero if and only if $Y,\,S$ are independent.
@@ 4 Inductive Biases of DP-based Fair Supervised Learning
subject $\text{\rm subject to}\;\;\;\rho\bigl{(}\widehat{Y},S\bigr{)}\leq\epsilon$ (2)
@@
${TV}(P_{Y},Q_{Y})\,\mathrel{\mathop{:}}=\,\frac{1}{2}\sum_{y\in\mathcal{Y}}\,\bigl{|}P_{Y}(y)-Q_{Y}(y)\bigr{|}$
@@
$\forall s\in\mathcal{S}\mathrel{\mathop{:}}\quad\mathrm{TV}\Bigl{(}P_{\,\widehat{Y}|S=s},P_{\,{Y}|S=s_{\max}}\Bigr{)}\,\leq\,\epsilon\bigl{(}1+\frac{1}{2\delta}\bigr{)}$
@@
$\forall s\in\mathcal{S}\mathrel{\mathop{:}}\;\;P_{\widehat{Y}|S=s}\,=\,P_{{Y}|S=s_{\max}}$
@@
    $\mathbb{E}_{s\sim P_{S}}\Bigl{[}\mathrm{TV}\Bigl{(}P_{\widehat{Y}|S=s},P_{{Y}|S=s_{\max}}\Bigr{)}\Bigr{]}\leq\bigl{(}1+\frac{1}{2\delta}\bigr{)}\sqrt{\frac{2\epsilon}{\log e}}$
@@
    $\mathbb{E}_{s\sim P_{S}}\Bigl{[}\mathrm{TV}\Bigl{(}P_{\widehat{Y}|S=s},P_{{Y}|S=s_{\max}}\Bigr{)}\Bigr{]}\leq\bigl{(}1+\frac{1}{2\delta}\bigr{)}u(\epsilon)$
@@
    $\mathbb{E}_{s\sim P_{S}}\Bigl{[}\mathrm{TV}\Bigl{(}P_{\widehat{Y}|S=s},P_{{Y}|S=s_{\max}}\Bigr{)}\Bigr{]}\leq\bigl{(}1+\frac{1}{2\delta}\bigr{)}u(r\epsilon)$
@@
We remark the difference between the bias levels shown for the DDP case in Theorem 1 and the other dependence metrics in Theorem 2. The bias level for a DDP-based fair leaner could be considerably stronger than that of mutual information, ERMI, and maximal correlation-based fair learners, as the maximum the total variations in Theorem 1 is replaced by their expected value over $P_{S}$ in Theorem 2.
@@ 4.1 Extending the Theoretical Results to Randomized Prediction Rules
In this subsection, we also consider the possibility of a randomized mapping from $(\mathbf{X},S)$ to $Y$. Such a possibility needs to be considered when the actual label $Y$ may not be deterministically determined by $\mathbf{X},S$. Therefore, we formulate and analyze the following generalization of the problem formulation in (1) where we attempt to find the conditional distribution $P_{Y|\mathbf{X},S}$:
@@
subject $\text{\rm subject to}\;\;\;\rho\bigl{(}\widehat{Y},S\bigr{)}\leq\epsilon$ (4)
@@
$\ell_{\mathrm{TV}}(P,Q)\,=\,\min_{\begin{subarray}{c}M_{\widehat{Y},Y}\mathrel{\mathop{:}}\>M_{\widehat{Y}}=P\\ \qquad\;\;\,M_{Y}=Q\end{subarray}}\;\;\mathbb{E}_{M}\Bigl{[}\ell_{0/1}\bigl{(}\widehat{Y},Y\bigr{)}\Bigr{]}.$
Therefore, if under both $Q$ and $P$, $Y$ is determined deterministically by $\mathbf{X},S$, the above TV-loss will be the same as the expected 0/1 loss of the deterministic classification rule following such $Q_{\widehat{Y}|\mathbf{X},S}$.
In the following, we attempt to relax the assumptions in Theorems 1-2 to apply them to learning settings where $Y$ may not be completely determined by $\mathbf{X},S$, and instead we use the following assumption on the distribution $p_{\mathbf{X},Y,S}$.
Assumption 1.
Suppose that the random vector $\mathbf{X}$ can be decomposed to $\bigl{[}g(S),\widetilde{\mathbf{X}}\bigr{]}$, where $g$ is a deterministic function, and under the true distribution $p_{\mathbf{X},Y,S}$, $\widetilde{\mathbf{X}}$ satisfies $\widetilde{\mathbf{X}}\bot S$, i.e., is independent from $S$, and $\widetilde{\mathbf{X}}\bot S\,\big{|}\,Y$, i.e, $\widetilde{\mathbf{X}}$ is conditionally independent from $S$ given $Y$.
Our next theorem shows the conclusions in the previous theorems further hold under Assumption 1.
@@
Consider the settings in Theorem 1 and Theorem 2 where we instead consider the generalized formulation (3) and do not require that $Y$ is a function of $\mathbf{X},S$. If Assumption 1 holds, then the conclusions in Theorems 1,2 will remain ture.
@@
$\forall\mathbf{x}\in\mathcal{X},y\in\mathcal{Y},s\in\mathcal{S}\mathrel{\mathop{:}}\quad\frac{P\bigl{(}\mathbf{x}\,\big{|}\,y,s\bigr{)}}{P\bigl{(}\mathbf{x}\,\big{|}\,s\bigr{)}}\,=\,\phi\bigl{(}\mathbf{x},y\bigr{)}.$
Note that this assumption is equivalent to a $s$-independent ratio $\frac{P(\mathbf{x}|y,s)}{P(\mathbf{x}|s)}$.

### v2

@@ Abstract
Fair supervised learning algorithms assigning labels with little dependence on a sensitive attribute have attracted great attention in the machine learning community. While the demographic parity (DP) notion has been frequently used to measure a model’s fairness in training fair classifiers, several studies in the literature suggest potential impacts of enforcing DP in fair learning algorithms. In this work, we analytically study the effect of standard DP-based regularization methods on the conditional distribution of the predicted label given the sensitive attribute. Our analysis shows that an imbalanced training dataset with a non-uniform distribution of the sensitive attribute could lead to a classification rule biased toward the sensitive attribute outcome holding the majority of training data. To control such inductive biases in DP-based fair learning, we propose a sensitive attribute-based distributionally robust optimization (SA-DRO) method improving robustness against the marginal distribution of the sensitive attribute. Finally, we present several numerical results on the application of DP-based learning methods to standard centralized and distributed learning problems. The empirical findings support our theoretical results on the inductive biases in DP-based fair learning algorithms and the debiasing effects of the proposed SA-DRO method. The project code is available at github.com/lh218/Fairness-IB.git.
@@ 1 Introduction
A widely-used approach to fair supervised learning is to include a fairness regularization penalty term in the learning objective that quantifies the level of fairness violation according to a fairness notion. A standard fairness notion is the demographic parity (DP) aiming toward a statistically independent prediction variable $\widehat{Y}$ of a sensitive attribute $S$. Therefore, a DP-based fairness regularization metric should be a measure of the dependence of the prediction $\widehat{Y}$ on the sensitive attribute $S$. In the literature, several dependence measures from statistics and information theory have been attempted to develop DP-based fair learning methodologies [Zafar et al., 2017, Mary et al., 2019, Baharlouei et al., 2019, Rezaei et al., 2020, Cho et al., 2020a, b, Lowy et al., 2022].
In practice, the applications of standard DP-based fair classification methods usually succeed in significantly reducing the DP fairness violation, while the model’s original accuracy on test data can be mostly preserved. Therefore, an accuracy-based evaluation of the DP-based trained models often suggests that the improvement in the DP fairness metric can be significantly higher than the loss in the model’s prediction accuracy. On the other hand, well-known studies including [Dwork et al., 2012] and [Hardt et al., 2016] have raised concerns about the potential impacts of DP-based fairness evaluation, which can disproportionately increase the inaccuracy rate among minority subgroups. To address the concerns, Hardt et al. [2016] propose and promote a different fairness notion, equalized odds (EO), where the goal is a prediction variable $\widehat{Y}$ that is conditionally independent of the sensitive attribute $S$ given the true label $Y$.
@@
We present the results of several numerical experiments on the potential biases of DP-based fair classification methodologies to the sensitive attribute possessing the majority in the dataset. Our empirical findings are consistent with the theoretical results, suggesting the inductive biases of DP-based fair classification rules toward the sensitive attribute-based majority group. On the other hand, our results indicate that the SA-DRO-based fair learning method results in fair classification rules with a lower bias toward the label distribution under the majority sensitive attribute. The following is a summary of this work’s main contributions:
@@ 2 Related Works
Fairness Violation Metrics. In this work, we focus on the learning frameworks aiming toward demographic parity (DP). Since enforcing DP to strictly hold could be costly and damaging to the learner’s performance, the machine learning literature has proposed applying several metrics assessing the dependence between random variables, including: the mutual information: [Kamishima et al., 2011, Rezaei et al., 2020, Zhang et al., 2018, Cho et al., 2020a, Roh et al., 2020], Pearson correlation [Zafar et al., 2017, Beutel et al., 2019], kernel-based maximum mean discrepancy: [Prost et al., 2019], kernel density estimation of the difference of demographic parity (DDP) measures [Cho et al., 2020b], the maximal correlation [Mary et al., 2019, Baharlouei et al., 2019, Grari et al., 2019, 2021], and the exponential Renyi mutual information [Lowy et al., 2022]. In our analysis, we mostly focus on a DP-based fair regularization scheme, while we show only weaker versions of the inductive biases could further hold in the case of mutual information and maximal correlation-based fair learning algorithms.
In addition to DP, the notions of equalized odds and equal opportunity [Hardt et al., 2016] are standard fairness notions in the literature, where the learner aims for a conditionally independent decision variable $\widehat{Y}$ of sensitive attribute $S$ given label $Y$. We note that the mentioned frameworks based on dependence measures can be aimed at equalized odds, where the dependence measure should be conditioned to label $Y$. Hence, our findings do not apply to the equalized odds fairness notion and the extension of the dependence measure-based learning algorithms aiming equalized odds.
Fair Classification Algorithms. Fair machine learning algorithms can be classified into three main categories: pre-processing, post-processing, and in-processing. Pre-processing algorithms [Feldman et al., 2015, Zemel et al., 2013, Calmon et al., 2017] transform biased data features into a new space where labels and sensitive attributes are statistically independent. Post-processing methods such as [Hardt et al., 2016, Pleiss et al., 2017] aim to alleviate the discriminatory impact of a classifier by modifying its ultimate decision. The focus of our work focus is only on in-processing approaches regularizing the training process toward DP-based fair models. Also, [Hashimoto et al., 2018, Wang et al., 2020, Lahoti et al., 2020] propose distributionally robust optimization (DRO) for fair classification; however, unlike our method, these works do not apply DRO on the sensitive attribute distribution to reduce the biases.
Fairness-aware Imbalanced Learning. To address the challenges of generalization in machine learning models, particularly when handling highly imbalanced classes and limited samples within each class, some well-known imbalanced learning methods like [Lin et al., 2017] and [Cao et al., 2019] have been proposed. More specifically, several articles [Iosifidis and Ntoutsi, 2020], [Subramanian et al., 2021], [Deng et al., 2022] and [Tarzanagh et al., 2023] extended to fairness-aware imbalanced learning dealing with imbalanced subgroups based on sensitive attributes. Compared to those methods, our SA-DRO method has more flexibility in exploring the accuracy-inductive bias trade-off controlled by varying the coefficient of the regularization term.
@@ 3.1 Fair Supervised Learning
In the supervised learning problem, the learner selects a function $f\in\mathcal{F}$ where $\mathcal{F}$ is the set of prediction functions mapping the observed $(\mathbf{X},S)$ to the label space $\mathcal{Y}$. We use loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ to quantify the loss $\ell(y,\widehat{y})$ when predicting $\widehat{y}$ under a true label $y$. Specifically, we consider the 0/1 loss $\ell_{0/1}(\widehat{y},y)=\mathbf{1}(\widehat{y}\neq y)$, where $\mathbf{1}(\cdot)$ denotes the indicator function. The primary goal of the fair supervised learner is to find prediction rules $f\in\mathcal{F}$ achieving smaller values of risk function $\mathbb{E}_{(\mathbf{X},Y,S)\sim P}\bigl{[}\ell(f(\mathbf{X},S),Y)\bigr{]}$ while having little dependence on $S$ according to the factors explained in the next subsections.
@@ 3.2 Fairness Criteria
  • Equalized Odds (EO) [Hardt et al., 2016] is a fairness condition requiring the predicted label $Y$ to be conditionally independent from sensitive attribute $S$ given actual label $Y$, i.e. for every $s\in\mathcal{S},y,\widehat{y}\in\mathcal{Y}$
@@
    $\centering\mathrm{DEO}(\widehat{Y},S|Y)=\sum_{s\in\mathcal{S},y,\hat{y}\in\mathcal{Y}}\Bigl{|}\,\@add@centering$ $P(\widehat{Y}=\widehat{y}\,\big{|}\,Y=y,S=s)$
    $-P\bigl{(}\widehat{Y}=\widehat{y}\,\big{|}\,Y=y\bigr{)}\Bigr{|}$
@@ 3.3 Dependence Measures for Fair Supervised Learning
  • Mutual Information (MI): Mutual information $I(Y;S)$ is a standard measure of the dependence between random variables $Y$ and $S$ used for developing fair learning methods [Cho et al., 2020a]. The mutual information $I(Y;S)$ is defined as
    $I(Y;S):=\sum_{y\in\mathcal{Y},s\in\mathcal{S}}P_{Y,S}(y,s)\log\frac{P_{Y,S}(y,s)}{P_{Y}(y)P_{S}(s)}$
@@
  • Maximal Correlation (MC): The maximal correlation $\rho_{m}(Y,S)$ is the maximum Pearson correlation $\rho_{P}\bigl{(}f(Y),g(S)\bigr{)}=\frac{\mathrm{Cov}(f(Y),g(S))}{\sqrt{\mathrm{Var}(f(Y))\mathrm{Var}(g(S))}}$ between $f(Y)$ and $g(S)$ over all functions $f,\,g$. The maximal correlation can be simplified to the optimal value of the following optimization:
    $\rho_{m}(Y,S):=\sup_{\begin{subarray}{c}f,g:\;\mathbb{E}[f(Y)]=\mathbb{E}[g(S)]=0\\ \qquad\,\mathbb{E}[f^{2}(Y)]=\mathbb{E}[g^{2}(S)]=1\end{subarray}}\mathbb{E}\bigl{[}f(Y)g(S)\bigr{]}$
    Maximal correlation has been utilized as a measure of demographic parity in the literature on fair learning algorithms [Mary et al., 2019, Baharlouei et al., 2019].
  • Exponential Rényi Mutual Information (ERMI): The ERMI between random variables $Y$ and $S$, which is considered by Lowy et al. [2022] as the dependence measure of fairness penalty, is
    $\rho_{E}(Y,S):=\chi^{2}(P_{Y,S};P_{Y}\times P_{S}),$
    i.e, the $\chi^{2}$-divergence between the joint distribution $P_{Y,S}$ and the product of marginal distributions $P_{Y}\times P_{S}$. Similar to KL-divergence, $\chi^{2}$-divergence is an $f$-divergence $d_{f}(P,Q)$ with $f(t)=(t-1)^{2}$. Similar to the previous two dependence measures, $\rho_{E}(Y,S)=0$ if and only if $Y,\,S$ are independent.
@@ 4 Inductive Biases of DP-based Fair Supervised Learning
subject $\text{\rm subject to}\;\;\;\rho\bigl{(}\widehat{Y},S\bigr{)}\leq\epsilon$
@@
${TV}(P_{Y},Q_{Y})\,:=\,\frac{1}{2}\sum_{y\in\mathcal{Y}}\,\bigl{|}P_{Y}(y)-Q_{Y}(y)\bigr{|}$
@@
$\forall s\in\mathcal{S}:\quad\mathrm{TV}\Bigl{(}P_{\,\widehat{Y}|S=s},P_{\,{Y}|S=s_{\max}}\Bigr{)}\,\leq\,\bigl{(}\frac{1}{2}+\frac{1}{4\delta}\bigr{)}\epsilon$
@@
$\forall s\in\mathcal{S}:\;\;P_{\widehat{Y}|S=s}\,=\,P_{{Y}|S=s_{\max}}$
@@
    $\mathbb{E}_{s\sim P_{S}}\Bigl{[}\mathrm{TV}\Bigl{(}P_{\widehat{Y}|S=s},P_{{Y}|S=s_{\max}}\Bigr{)}\Bigr{]}\leq\bigl{(}\frac{1}{2}+\frac{1}{4\delta}\bigr{)}\sqrt{\frac{2\epsilon}{\log e}}$
@@
    $\mathbb{E}_{s\sim P_{S}}\Bigl{[}\mathrm{TV}\Bigl{(}P_{\widehat{Y}|S=s},P_{{Y}|S=s_{\max}}\Bigr{)}\Bigr{]}\leq\bigl{(}\frac{1}{2}+\frac{1}{4\delta}\bigr{)}u(\epsilon)$
@@
    $\mathbb{E}_{s\sim P_{S}}\Bigl{[}\mathrm{TV}\Bigl{(}P_{\widehat{Y}|S=s},P_{{Y}|S=s_{\max}}\Bigr{)}\Bigr{]}\leq\bigl{(}\frac{1}{2}+\frac{1}{4\delta}\bigr{)}u(r\epsilon)$
@@
We remark the difference between the bias levels shown for the DDP case in Theorem 1 and the other dependence metrics in Theorem 2. The bias level for a DDP-based fair learner could be considerably stronger than that of mutual information, ERMI, and maximal correlation-based fair learners, as the wort-case of total variations in Theorem 1 is replaced by their expectation according to $P_{S}$ in Theorem 2.
@@ 4.1 Extending the Theoretical Results to Randomized Prediction Rules
Here, we consider the possibility of a randomized mapping from $(\mathbf{X},S)$ to $Y$. Such a possibility needs to be considered when the actual label $Y$ may not be deterministically determined by $\mathbf{X},S$. Therefore, we formulate and analyze the following generalization of the problem formulation in (1) where we attempt to find the conditional distribution $P_{Y|\mathbf{X},S}$:
@@
subject $\text{\rm subject to}\;\;\;\rho\bigl{(}\widehat{Y},S\bigr{)}\leq\epsilon$
@@
$\ell_{\mathrm{TV}}(P,Q)\,=\,\min_{\begin{subarray}{c}M_{\widehat{Y},Y}:\>M_{\widehat{Y}}=P\\ \qquad\;\;\,M_{Y}=Q\end{subarray}}\;\;\mathbb{E}_{M}\Bigl{[}\ell_{0/1}\bigl{(}\widehat{Y},Y\bigr{)}\Bigr{]}.$
Therefore, if under both $Q$ and $P$, $Y$ is determined deterministically by $\mathbf{X},S$, the above TV-loss will be the same as the expected 0/1 loss of the deterministic classification rule following such $Q_{\widehat{Y}|\mathbf{X},S}$. In the following theorem, we attempt to relax the assumptions in Theorems 1-2 to apply them to learning settings where $Y$ may not be completely determined by $\mathbf{X},S$.
@@
Consider the settings in Theorem 1 and Theorem 2 where we instead consider the generalized formulation (2) and do not require that $Y$ is a function of $\mathbf{X},S$. Suppose a function $\phi:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ exists such that the underlying distribution $P_{\mathbf{X},Y,S}$ satisfies the following property on the ratio between conditional distributions $P_{\mathbf{X}|Y,S}$ and $P_{\mathbf{X}|S}$:
$\forall\mathbf{x}\in\mathcal{X},y\in\mathcal{Y},s\in\mathcal{S}:\quad\frac{P\bigl{(}\mathbf{x}\,\big{|}\,y,s\bigr{)}}{P\bigl{(}\mathbf{x}\,\big{|}\,s\bigr{)}}\,=\,\phi\bigl{(}\mathbf{x},y\bigr{)}.$ (3)
Then, the conclusions in Theorems 1,2 will remain valid.
@@
Note that the assumption in the above theorem is equivalent to a $s$-independent ratio $\frac{P(\mathbf{x}|y,s)}{P(\mathbf{x}|s)}$. In particular, this assumption will hold if the random vector $\mathbf{X}$ can be decomposed to $\bigl{[}g(S),\widetilde{\mathbf{X}}\bigr{]}$, where $g$ is a deterministic function, and under the true distribution $p_{\mathbf{X},Y,S}$, $\widetilde{\mathbf{X}}$ satisfies $\widetilde{\mathbf{X}}\bot S$, i.e., is independent from $S$, and $\widetilde{\mathbf{X}}\bot S\,\big{|}\,Y$, i.e, $\widetilde{\mathbf{X}}$ is conditionally independent from $S$ given $Y$.
Finally, we attempt to further relax the assumption in Theorem 3 when the distribution ratio $P(\mathbf{x}|y,s)/P(\mathbf{x}|s)$ may not be completely independent of the outcome $S=s$. The next theorem shows a quantification of the deviation from the assumption and how much it can impact the result.
Theorem 4.
Consider the setting of Theorem 3 and the formulation (2). We consider the TV-based dependence $\rho_{TV}(Y,S):=\mathbb{E}_{s\sim P_{S}}\bigl{[}\mathrm{TV}(P_{Y|S=s},P_{Y})\bigr{]}$ in the problem. Suppose for functions $\phi_{L},\,\phi_{U}:\,\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, the following holds for every $\mathbf{x}\in\mathcal{X},y\in\mathcal{Y},s\in\mathcal{S}$:
$\phi_{L}(\mathbf{x},y)\,\leq\,\frac{p\bigl{(}\mathbf{x}|y,s\bigr{)}}{p\bigl{(}\mathbf{x}|s\bigr{)}}\,\leq\,\phi_{U}(\mathbf{x},y).$
Define $\Delta(\mathbf{x},y)=\phi_{U}(\mathbf{x},y)-\phi_{L}(\mathbf{x},y)$. Then, if $\frac{\epsilon}{2}\geq\mathbb{E}_{P_{X}P_{Y|S=s_{\max}}}\bigl{[}\Delta(\mathbf{x},y)\bigr{]}$, for the optimal $Q^{*}_{\widehat{Y}|\mathbf{X},S}$, $P_{\widehat{Y},\mathbf{X},S}=Q^{*}_{\widehat{Y}|\mathbf{X},S}\cdot P_{\mathbf{X},S}$ satisfies
@@
Proof.
We defer the proof to the Appendix. ∎